OpenStudy (anonymous):

differentiate the function f(x_=sin(cos(x^2))

OpenStudy (anonymous):

[cos(cos(x^2)] [-sin(x^2)(2x)]

OpenStudy (anonymous):

\[f(x)= \sin(\cos(x^{2})) \]\[f' = -2xsin(x^{2})(\cos(x^{2}) = -4(\cos^{2}(x^{2})-1)\]